# FedP2EFT: Federated Learning to Personalize PEFT for Multilingual LLMs

### v2

@@ A.3 Complementary Approaches and Base Models
Standard FL. Standard FL involves training a single global model. Given a pretrained LLM, we run FedAvg on the seen pool of clients, where $10\%$ of participating clients are sampled every round to train the model before sending the weights back for aggregation. In our XNLI and MasakhaNEWS setup, we do full fine-tuning of mBERT, setting each client’s learning rate to 5e55e-5 and running FedAvg for 100 rounds. In our Fed-Aya setup, we adopt the training recipe from FedLLM-Bench (Ye et al. 2024) for MobileLLaMA-1.4B, where we do PEFT with LoRA applied to query and value attention weights ($r=16$, $\alpha_{lora}=32$, dropout=0.05=0.05) for 200 rounds. We use the cosine learning rate decay over 200 rounds with initial learning rate $2e^{-5}$ and minimum learning rate $1e^{-6}$.
Personalized FL. We train personalized base models using FedDPA-T (Yang et al. 2024b) and DEPT(SPEC) (Iacob et al. 2025) in our XNLI setup. FedDPA-T proposed having two separate LoRA adapters, one of which is shared (global) and the other is kept locally for each client (local). We adopted their training recipe for sequence classification, where the classifier is shared together with the global LoRA modules and the local LoRA modules stay local. LoRA modules are only applied to query and value attention weights (r=8, $\alpha_{lora}=8$, dropout=$0.05$). We set the learning rate to $5e^{-5}$.
@@ A.4 Baselines and FedP2EFT
LoRA PEFT (Hu et al. 2022). We search for the learning rate $1e^{-5},1e^{-4},1e^{-3}]$ and the number of epochs $[1,2,3]$ and find that the learning rate $1e^{-4}$ with $2$ epochs had the best performance on the train set. We fixed $\alpha_{lora}=2r$. To ensure a similar inference budget across baselines, we set the number of epochs to $2$ for all our experiments.
@@
FedL2P (Lee et al. 2023). As FedL2P requires a validation set for outer-loop bi-level optimization and federated early stopping, we split the train set of every client $80\%$ train and $20\%$ validation. Following FedL2P, we set the federated early stopping patience to 50 rounds, MLP hidden dimension is set to 100, the inner-loop learning rate to be the same as finetuning, $1e^{-4}$, and the hypergradient hyperparameters, $Q=3,\psi=0.1$ with hypergradient clamping of [1,1][-1,1]. We use Adam for both the inner-loop and outer-loop optimizers and search for the learning rate for the MLP (LRNet) $1e^{-5},1e^{-4},1e^{-3}]$ and the learnable post-multiplier learning rate $\tilde{\eta}$ $1e^{-4},1e^{-5},1e^{-6}]$, picking $1e^{-4}$ and $1e^{-6}$ to be the best respectively. Finally, we use $3$ outer-loop steps with an effective outer-loop batch size of $16$.
FedP2EFT (Section 3.3). We set $\alpha_{r\_mul}=2$, our resulting $r_{init}$ is, hence, $32$ since our $r_{\text{max target}}=16$ for all experiments. Following our standard LoRA fine-tuning baseline, we adopt the same learning rate and $\alpha_{lora}$, 1e41e-4 and $2r_{init}$ respectively. The learning rate of $\bm{\lambda}$, on the other hand, is searched $1e^{-1},1e^{-2},1e^{-3},1e^{-4}]$, and we pick $1e^{-2}$ for all experiments. All $\lambda$ values are initialized to $1e^{-4}$. The MLP hidden dimension is set to $2\times$ the input dimension, which is model dependent. We clamp the output of the MLP as well as $\lambda$ with a minimum value of 1e41e-4 in the forward pass during training. Following FedL2P, we use a straight-through estimator (Bengio, Léonard, and Courville 2013) after clamping to propagate gradients. We initialize the weights of MLP with Xavier initialization (Glorot and Bengio 2010) using the normal distribution with a gain value of 1e61e-6 and the bias with a constant 1e41e-4. Lastly, we set $\alpha_{s}=1e^{+2}$ and $\alpha_{p}=1e^{-2}$ for all experiments.
@@ A.5 Experiments Compute Resource
We used Flower (Beutel et al. 2020) [^8] https://github.com/adap/flower, which uses Ray (Moritz et al. 2018) [^9] https://github.com/ray-project/ray for distributed training and inference, for all our experiments. Each experiment in our XNLI and MasakhaNEWS setup was run with 4 to 6 NVIDIA A40s GPUs, with each client running on 1 GPU at a time. Each FedP2EFT federated training run took less than an hour in total while FedL2P training run took a few hours for each rank budget. As accuracy is cheap to compute, each inference run only took a few minutes to evaluate on all clients. Our Fed-Aya setup with Llama-3.2-3B, on the other hand, took under a day to federatedly train FedP2EFT and 1-3 days to train FedL2P for each rank budget on 8 NVIDIA A100s 80GB. Due to the higher cost of auto-regressive generative runs compared to classification, each inference run took roughly 1-3 hours.
@@ Appendix C On BayesTune
BayesTune (Kim and Hospedales 2023) is a sparse fine-tuning approach that leverages a Bayesian framework to identify and update only the most important parameters of the model. It introduces a per-parameter scaling coefficient, where the magnitude of this coefficient governs the degree to which the corresponding parameter is updated during fine-tuning. These scaling coefficients, along with the model parameters, are explored and sampled from their posterior distribution using SGLD. To encourage sparsity, the Laplace and Gamma priors are imposed on the model parameters and their scaling coefficients respectively as discussed in Section 3.2.
@@
However, as mentioned in Section 3.2, directly using BayesTune in our approach will not only incur an intractable cost but also be difficult to optimize as the output dimension of our MLP will be too big and the likelihood of getting stuck in a suboptimal local minimum is high. This limitation can also be observed in existing FL hypernetwork approaches (Shamsian et al. 2021) where the experiments were constrained to learning the personalized parameters of a tiny LeNet-5. Similarly, the authors of FedL2P (Lee et al. 2023) found that extending these hypernetwork approaches to ResNet-18 is costly and often struggle to converge. To make federated hyperparameter optimization feasible, we scale down this output space by learning optimal LoRA rank coefficients instead. While doing so improves convergence stability, the output dimension scales linearly with the number of LoRA ranks, thereby also susceptible to the aforementioned limitation at higher ranks.
@@ Appendix E Additional Results
Lang LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Additional Quantitative Results. We show extended results for 1) our MasakhaNEWS setup in Tables 7 and 8, 2) our XNLI setup with FedDPA-T used to pretrain the base model in Table 9, and 3) our Fed-Aya setup with off-the-shelf Llama3 in Tables 10 and 11 for seen and unseen clients respectively.
We also present additional results to supplement the results in the main paper, namely 1) our XNLI setup with Standard FL (full fine-tuning) used to pretrain the base model in Tables 13 and 14, 2) our XNLI setup to show that FedP2EFT is complementary with other pFL methods in Table 12, and 3) our Fed-Aya setup with MobileLLaMA as the base model in Tables 15 and 16.
@@
Approach LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Approach LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Table 15: Average METEOR/ROUGE-1/ROUGE-L of each language for seen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (Ye et al. 2024) and the resulting base model is personalized to each client given a baseline approach.
@@
Table 16: Average METEOR/ROUGE-1/ROUGE-L of each language for unseen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (Ye et al. 2024) and the resulting base model is personalized to each client given a baseline approach.

### v3

@@ A.3 Complementary Approaches and Base Models
Standard FL. Standard FL involves training a single global model. Given a pretrained LLM, we run FedAvg on the seen pool of clients, where $10\%$ of participating clients are sampled every round to train the model before sending the weights back for aggregation. In our XNLI and MasakhaNEWS setup, we do full fine-tuning of mBERT, setting each client’s learning rate to 5e55e-5 and running FedAvg for 100 rounds. In our Fed-Aya setup, we adopt the training recipe from FedLLM-Bench (fedllm-bench) for MobileLLaMA-1.4B, where we do PEFT with LoRA applied to query and value attention weights ($r=16$, $\alpha_{lora}=32$, dropout=0.05=0.05) for 200 rounds. We use the cosine learning rate decay over 200 rounds with initial learning rate $2e^{-5}$ and minimum learning rate $1e^{-6}$.
Personalized FL. We train personalized base models using FedDPA-T (FedDPA) and DEPT(SPEC) (DEPT) in our XNLI setup. FedDPA-T proposed having two separate LoRA adapters, one of which is shared (global) and the other is kept locally for each client (local). We adopted their training recipe for sequence classification, where the classifier is shared together with the global LoRA modules and the local LoRA modules stay local. LoRA modules are only applied to query and value attention weights (r=8, $\alpha_{lora}=8$, dropout=$0.05$). We set the learning rate to $5e^{-5}$.
@@ A.4 Baselines and FedP2EFT
LoRA PEFT (hu2021lora). We search for the learning rate $1e^{-5},1e^{-4},1e^{-3}]$ and the number of epochs $[1,2,3]$ and find that the learning rate $1e^{-4}$ with $2$ epochs had the best performance on the train set. We fixed $\alpha_{lora}=2r$. To ensure a similar inference budget across baselines, we set the number of epochs to $2$ for all our experiments.
@@
FedL2P (royson2023fedl2p). As FedL2P requires a validation set for outer-loop bi-level optimization and federated early stopping, we split the train set of every client $80\%$ train and $20\%$ validation. Following FedL2P, we set the federated early stopping patience to 50 rounds, MLP hidden dimension is set to 100, the inner-loop learning rate to be the same as finetuning, $1e^{-4}$, and the hypergradient hyperparameters, $Q=3,\psi=0.1$ with hypergradient clamping of [1,1][-1,1]. We use Adam for both the inner-loop and outer-loop optimizers and search for the learning rate for the MLP (LRNet) $1e^{-5},1e^{-4},1e^{-3}]$ and the learnable post-multiplier learning rate $\tilde{\eta}$ $1e^{-4},1e^{-5},1e^{-6}]$, picking $1e^{-4}$ and $1e^{-6}$ to be the best respectively. Finally, we use $3$ outer-loop steps with an effective outer-loop batch size of $16$.
FedP2EFT (Section 3.3). We set $\alpha_{r\_mul}=2$, our resulting $r_{init}$ is, hence, $32$ since our $r_{\text{max target}}=16$ for all experiments. Following our standard LoRA fine-tuning baseline, we adopt the same learning rate and $\alpha_{lora}$, 1e41e-4 and $2r_{init}$ respectively. The learning rate of $\bm{\lambda}$, on the other hand, is searched $1e^{-1},1e^{-2},1e^{-3},1e^{-4}]$, and we pick $1e^{-2}$ for all experiments. All $\lambda$ values are initialized to $1e^{-4}$. The MLP hidden dimension is set to $2\times$ the input dimension, which is model dependent. We clamp the output of the MLP as well as $\lambda$ with a minimum value of 1e41e-4 in the forward pass during training. Following FedL2P, we use a straight-through estimator after clamping to propagate gradients. We initialize the weights of MLP with Xavier initialization using the normal distribution with a gain value of 1e61e-6 and the bias with a constant 1e41e-4. Lastly, we set $\alpha_{s}=1e^{+2}$ and $\alpha_{p}=1e^{-2}$ for all experiments.
@@ A.5 Experiments Compute Resource
We used Flower (beutel2020flower) [^9] https://github.com/adap/flower, which uses Ray (moritz2018ray) [^10] https://github.com/ray-project/ray for distributed training and inference, for all our experiments. Each experiment in our XNLI and MasakhaNEWS setup was run with 4 to 6 NVIDIA A40s GPUs, with each client running on 1 GPU at a time. Each FedP2EFT federated training run took less than an hour in total while FedL2P training run took a few hours for each rank budget. As accuracy is cheap to compute, each inference run only took a few minutes to evaluate on all clients. Our Fed-Aya setup with Llama-3.2-3B, on the other hand, took under a day to federatedly train FedP2EFT and 1-3 days to train FedL2P for each rank budget on 8 NVIDIA A100s 80GB. Due to the higher cost of auto-regressive generative runs compared to classification, each inference run took roughly 1-3 hours.
@@ Appendix C On BayesTune
BayesTune (kim2023bayestune) is a sparse fine-tuning approach that leverages a Bayesian framework to identify and update only the most important parameters of the model. It introduces a per-parameter scaling coefficient, where the magnitude of this coefficient governs the degree to which the corresponding parameter is updated during fine-tuning. These scaling coefficients, along with the model parameters, are explored and sampled from their posterior distribution using SGLD. To encourage sparsity, the Laplace and Gamma priors are imposed on the model parameters and their scaling coefficients respectively as discussed in Section 3.2.
@@
However, as mentioned in Section 3.2, directly using BayesTune in our approach will not only incur an intractable cost but also be difficult to optimize as the output dimension of our MLP will be too big and the likelihood of getting stuck in a suboptimal local minimum is high. This limitation can also be observed in existing FL hypernetwork approaches (pfedhn) where the experiments were constrained to learning the personalized parameters of a tiny LeNet-5. Similarly, the authors of FedL2P (royson2023fedl2p) found that extending these hypernetwork approaches to ResNet-18 is costly and often struggle to converge. To make federated hyperparameter optimization feasible, we scale down this output space by learning optimal LoRA rank coefficients instead. While doing so improves convergence stability, the output dimension scales linearly with the number of LoRA ranks, thereby also susceptible to the aforementioned limitation at higher ranks.
@@ Appendix E Additional Results
Predominant Language Bias. A limitation of FedP2EFT, as discussed in the main paper, is that $\bm{\lambda}$ is heavily influenced by the client’s predominant language. To quantify this bias, we correlated each client’s predominant language proportion with the disparity in performance gain between predominant and minority languages in our Llama-3 3B setup, finding moderate positive Pearson coefficients of 0.6 (ROUGE-1), 0.56 (ROUGE-L), and 0.52 (METEOR). Potential future directions to address this include extracting per-language (rather than per-client) metadata as input to the PSG, or exploring multi-adapter approaches where each language receives a dedicated adapter.
@@
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
$\mathbf{r}$ 2
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.24/0.15/0.14 0.24/0.15/0.15 0.24/0.14/0.14 0.25/0.16/0.15
ar 0.34/0.07/0.07 0.33/0.07/0.07 0.32/0.06/0.06 0.34/0.07/0.07
es 0.38/0.42/0.39 0.39/0.42/0.40 0.37/0.41/0.38 0.40/0.43/0.40
en 0.36/0.41/0.34 0.37/0.43/0.35 0.32/0.33/0.27 0.40/0.45/0.38
fr 0.30/0.37/0.31 0.32/0.39/0.32 0.29/0.36/0.30 0.33/0.40/0.33
zh 0.10/0.13/0.13 0.11/0.13/0.13 0.10/0.13/0.12 0.10/0.13/0.12
pt 0.41/0.48/0.44 0.40/0.47/0.43 0.39/0.46/0.42 0.41/0.48/0.44
$\mathbf{r}$ 4
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.24/0.15/0.15 0.24/0.15/0.15 0.25/0.15/0.15 0.26/0.16/0.15
ar 0.35/0.07/0.07 0.33/0.07/0.07 0.33/0.07/0.07 0.36/0.07/0.07
es 0.40/0.43/0.40 0.39/0.42/0.39 0.38/0.42/0.39 0.40/0.44/0.40
en 0.38/0.43/0.36 0.36/0.41/0.34 0.34/0.37/0.30 0.41/0.47/0.39
fr 0.32/0.39/0.33 0.29/0.37/0.31 0.30/0.37/0.31 0.34/0.41/0.33
zh 0.11/0.14/0.14 0.11/0.13/0.13 0.11/0.13/0.13 0.11/0.13/0.13
pt 0.41/0.48/0.44 0.40/0.46/0.43 0.40/0.46/0.42 0.42/0.48/0.44
$\mathbf{r}$ 8
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.25/0.15/0.15 0.24/0.15/0.15 0.25/0.15/0.15 0.26/0.15/0.15
ar 0.35/0.07/0.07 0.32/0.07/0.07 0.35/0.07/0.07 0.37/0.08/0.08
es 0.40/0.43/0.40 0.38/0.41/0.39 0.38/0.42/0.39 0.40/0.43/0.40
en 0.39/0.45/0.37 0.36/0.40/0.33 0.36/0.41/0.34 0.41/0.47/0.39
fr 0.33/0.40/0.33 0.29/0.36/0.30 0.32/0.39/0.32 0.36/0.42/0.33
zh 0.11/0.14/0.14 0.11/0.13/0.13 0.10/0.13/0.13 0.12/0.13/0.13
pt 0.41/0.48/0.44 0.39/0.46/0.42 0.40/0.47/0.43 0.41/0.48/0.43
$\mathbf{r}$ 16
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.25/0.15/0.15 0.24/0.15/0.14 0.25/0.16/0.15 0.26/0.16/0.16
ar 0.36/0.08/0.07 0.32/0.06/0.06 0.36/0.07/0.07 0.37/0.08/0.08
es 0.41/0.44/0.41 0.38/0.41/0.38 0.40/0.43/0.40 0.42/0.45/0.41
en 0.40/0.46/0.38 0.34/0.37/0.30 0.38/0.43/0.36 0.41/0.47/0.39
fr 0.34/0.41/0.33 0.29/0.36/0.30 0.33/0.40/0.33 0.35/0.40/0.32
zh 0.12/0.14/0.13 0.11/0.12/0.12 0.11/0.14/0.13 0.11/0.13/0.13
pt 0.42/0.48/0.44 0.38/0.46/0.42 0.41/0.48/0.44 0.41/0.47/0.43
Table 12: Avg. METEOR/ROUGE-1/ROUGE-L for seen clients in our Fed-Aya setup. Base model is off-the-shelf Llama-3.1-8B-Instruct. FedL2P leads to out-of-memory exceptions due to its expensive second-order optimization.
$\mathbf{r}$ 2
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.17/0.12/0.11 0.17/0.12/0.11 0.15/0.12/0.12 0.16/0.11/0.10
ar 0.25/0.03/0.03 0.27/0.05/0.05 0.23/0.05/0.05 0.26/0.06/0.06
es 0.43/0.53/0.48 0.42/0.53/0.48 0.39/0.52/0.47 0.45/0.57/0.51
en 0.33/0.31/0.25 0.32/0.31/0.25 0.31/0.29/0.23 0.33/0.35/0.29
fr 0.44/0.50/0.50 0.21/0.56/0.56 0.53/0.61/0.61 0.53/0.60/0.60
zh 0.22/0.00/0.00 0.24/0.01/0.01 0.23/0.00/0.00 0.24/0.00/0.00
pt 0.37/0.44/0.42 0.33/0.43/0.41 0.31/0.36/0.34 0.35/0.45/0.44
ru 0.24/0.30/0.30 0.23/0.28/0.28 0.22/0.28/0.28 0.26/0.27/0.27
$\mathbf{r}$ 4
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.18/0.12/0.12 0.16/0.11/0.10 0.17/0.12/0.11 0.17/0.12/0.11
ar 0.27/0.04/0.04 0.25/0.03/0.03 0.24/0.05/0.05 0.26/0.06/0.06
es 0.47/0.58/0.53 0.41/0.52/0.47 0.40/0.53/0.47 0.50/0.59/0.54
en 0.34/0.34/0.28 0.31/0.31/0.25 0.30/0.29/0.23 0.31/0.35/0.30
fr 0.46/0.50/0.50 0.44/0.50/0.50 0.44/0.50/0.50 0.47/0.69/0.69
zh 0.23/0.01/0.01 0.24/0.01/0.01 0.23/0.00/0.00 0.26/0.00/0.00
pt 0.36/0.45/0.44 0.34/0.43/0.41 0.33/0.36/0.35 0.36/0.45/0.43
ru 0.25/0.34/0.34 0.26/0.35/0.35 0.22/0.27/0.27 0.21/0.17/0.17
$\mathbf{r}$ 8
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.17/0.10/0.09 0.16/0.09/0.09 0.17/0.13/0.12 0.18/0.10/0.09
ar 0.26/0.04/0.04 0.24/0.03/0.03 0.22/0.04/0.04 0.26/0.08/0.08
es 0.46/0.57/0.53 0.42/0.53/0.47 0.40/0.53/0.47 0.48/0.59/0.53
en 0.35/0.38/0.32 0.32/0.31/0.25 0.31/0.30/0.24 0.34/0.38/0.32
fr 0.46/0.50/0.50 0.44/0.50/0.50 0.44/0.50/0.50 0.29/0.43/0.43
zh 0.24/0.01/0.01 0.23/0.00/0.00 0.22/0.01/0.01 0.31/0.02/0.02
pt 0.36/0.45/0.44 0.31/0.40/0.39 0.34/0.43/0.41 0.37/0.46/0.44
ru 0.21/0.23/0.23 0.26/0.32/0.32 0.24/0.30/0.30 0.24/0.15/0.15
$\mathbf{r}$ 16
Lang. LoRA AdaLoRA BT-LoRA FedP2EFT
te 0.15/0.09/0.09 0.15/0.10/0.10 0.17/0.11/0.10 0.18/0.13/0.12
ar 0.27/0.06/0.06 0.23/0.04/0.04 0.27/0.05/0.05 0.24/0.08/0.08
es 0.46/0.55/0.51 0.43/0.54/0.48 0.41/0.53/0.48 0.42/0.53/0.48
en 0.33/0.38/0.31 0.31/0.30/0.24 0.31/0.30/0.25 0.35/0.39/0.31
fr 0.38/0.56/0.56 0.44/0.50/0.50 0.23/0.56/0.56 0.47/0.47/0.47
zh 0.25/0.01/0.01 0.24/0.00/0.00 0.23/0.01/0.01 0.33/0.04/0.04
pt 0.37/0.48/0.46 0.30/0.37/0.36 0.32/0.45/0.43 0.38/0.45/0.43
ru 0.25/0.28/0.28 0.27/0.35/0.35 0.24/0.30/0.30 0.29/0.22/0.22
Table 13: Avg. METEOR/ROUGE-1/ROUGE-L for unseen clients in our Fed-Aya setup. Base model is off-the-shelf Llama-3.1-8B-Instruct. FedL2P leads to out-of-memory exceptions due to its expensive second-order optimization.
@@
Additional Quantitative Results. We show extended results for 1) our MasakhaNEWS setup in Tables 7 and 8, 2) our XNLI setup with FedDPA-T used to pretrain the base model in Table 9, and 3) our Fed-Aya setup with off-the-shelf Llama3.2 3B in Tables 10 and 11 and off-the-shelf Llama3.1 8B in Tables 12 and 13 for seen and unseen clients respectively.
We also present additional results to supplement the results in the main paper, namely 1) our XNLI setup with Standard FL (full fine-tuning) used to pretrain the base model in Tables 15 and 16, 2) our XNLI setup to show that FedP2EFT is complementary with other pFL methods in Table 14, and 3) our Fed-Aya setup with MobileLLaMA as the base model in Tables 17 and 18.
@@
Table 17: Average METEOR/ROUGE-1/ROUGE-L of each language for seen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (fedllm-bench) and the resulting base model is personalized to each client given a baseline approach.
@@
Table 18: Average METEOR/ROUGE-1/ROUGE-L of each language for unseen clients of our Fed-Aya setup. The pretrained MobileLLaMA-1.4B model is trained using Standard FL with LoRA following FedLLM-Bench (fedllm-bench) and the resulting base model is personalized to each client given a baseline approach.